Identify whether the following compounds are a strong electrolyte, a weak electrolyte or a non-electrolyte. Hydrochloric...

  1. Identify whether the following compounds are a strong electrolyte, a weak electrolyte or a non-electrolyte.
    1. Hydrochloric acid
    2. Sodium iodide
    3. Sucrose
    4. Acetic acid
    5. Propanol

  1. Will mixing the following pairs of salts produce a precipitate? If so, what precipitate will form?
    1. KCl & Pb(NO3)2
    2. NaOH & HBr
    3. NH4NO3 & NaSO4
    4. CaCl2 & KOH
    5. LiNO3 & NaI

  1. A standard wine bottle has a volume of 750 mL. If wine has a volume percent of 11.0%, how many millilitres of ethanol (drinking alcohol) is inside that bottle?

  1. If 100 grams of sucrose (C12H22O11) was dissolved in 500 mL of solvent, what is the molarity of that solution?
  1. Using the solution made in question #4, if you wanted to make 250 mL of a diluted solution with a concentration of 0.100M, what volume of concentrated solution must you use?
  1. If you have 0.250M H3PO4 solution and you want to react it with 25.0 mL of 0.100M NaOH, how much H3PO4 solution must you use?

H3PO4 (aq) + 3NaOH (aq) → Na3PO4 (aq) + 3H2O (l)

  1. In regard to the reaction below, write an equilibrium constant expression and calculate the equilibrium constant if the equilibrium concentrations are: [H2S] = 1.25M, [S2] = 0.500M, [H2] = 0.100M.

2H2 (g) + S2 (g) → 2H2S(g)

  1. In regard to the equilibrium reaction below, what direction would the equilibrium shift if the following changes were made?

N2 (g) + 3H2 (g) → 2NH3 (g)

  1. Hydrogen gas was added
  2. Nitrogen was added
  3. Ammonia was removed
  4. Ammonia was added
  5. Both hydrogen gas and nitrogen were added

  1. What is solubility of silver sulfate (Ksp = 1.2 x 10-5)?
Answer #1
  1. Identify whether the following compounds are a strong electrolyte, a weak electrolyte or a non-electrolyte.
    1. Hydrochloric acid   = strong electrolyte
    2. Sodium iodide      = strong electrolyte
    3. Sucrose               = non- electrolyte
    4. Acetic acid          = weak electrolyte
    5. Propanol             = weak electrolyte

  1. Will mixing the following pairs of salts produce a precipitate? If so, what precipitate will form?
    1. KCl & Pb(NO3)2      = yes , PbCl2
    2. NaOH & HBr          = No
    3. NH4NO3 & NaSO4 = No
    4. CaCl2 & KOH         = yes , Ca(OH)2
    5. LiNO3 & NaI          = No

3)

volume % = 11.0 %

volume % = (volume of ethanol / total volume) x 100

11.0 = (volume of ethanol / 750 ) x 100

volume of ethanol = 82.5 mL

4)

mass of sucorse = 100 g

moles of sucrose = 100 / 342 = 0.292

Molarity = 0.292 / 0.500 = 0.585

Molarity = 0.585 M

5)

250 x 0.100 = 0.585 x V

V = 42.7 mL

volume of concentrated solution = 42.7 mL

6)

H3PO4 (aq) + 3NaOH (aq) → Na3PO4 (aq) + 3H2O (l)

mmoles of NaOH = 25.0 x 0.100 = 2.5

mmoles of H3PO4 = 2.5 / 3 = 0.833

0.250 x V = 0.833

V = 3.33

voume of H3PO4 = 3.33 mL

7)

  1. [H2S] = 1.25M, [S2] = 0.500M, [H2] = 0.100M.

2H2 (g) + S2 (g) → 2H2S(g)

Kc = [H2S]2 / [H2]2[S2]

    = (1.25)^2/ (0.100)^2(0.500)

Kc= 312.5